OpenStudy (anonymous):

Evaluate the limit ln lb+2/b+3l as b->infinity

OpenStudy (anonymous):

Those are absolute value bars around the numerator and denominator

OpenStudy (anonymous):

And I know it goes to 0, but I need to rewrite the numerator and denominator, and my algebra skills are failing me right now T_T

OpenStudy (anonymous):

Question for you: as b goes to infinity what is its relationship to 2 or 3?

OpenStudy (anonymous):

The 2 and 3 dont matter, so i cancelled them out

OpenStudy (anonymous):

good so now what would your fraction be?

OpenStudy (irishboy123):

think (b+2 + 1)/(b+3) - 1/(b+3)

OpenStudy (anonymous):

But then you get an indeterminate form in the ln

OpenStudy (anonymous):

because the fraction would become ln linfinity/infinityl

OpenStudy (anonymous):

thats not the case. what fraction would you have when you cancel out 2 and 3?

OpenStudy (anonymous):

ln lb/bl

OpenStudy (anonymous):

ok so if you just saw b/b what woulld you do with that?

OpenStudy (anonymous):

Declare it indeterminate and l'Hopital

OpenStudy (anonymous):

forget about calculus for a second. pretend you are in algebra class and your teacher said simplify the fraction

OpenStudy (tkhunny):

You have written "lb+2/b+3l " This is \(\left|b + \dfrac{2}{b} + 3\right|\). The limit does not exist as b increases without bound. Had you have written "l(b+2)/(b+3)l " This is \(\left|\dfrac{b+2}{b+3}\right|\). The limit does exist as b increases without bound. You do not need l'Hospiatl. The limit as b increases is 1. Divide Numerator and Denominator by b and show it to be so.

OpenStudy (anonymous):

Yes but if you plug in infinity Alex, its indeterminate

OpenStudy (anonymous):

Does that not matter?

OpenStudy (anonymous):

no no no...you dont plug in infinity until you have reduced it as much as possible

OpenStudy (anonymous):

Oh snap, I just realized how stupid I was

OpenStudy (anonymous):

b/b=1 T_T

OpenStudy (tkhunny):

There is no such thing as "plug in". Never do that. There is no such thing as "cancel". Never do that. Yes, it is an indeterminate form. Why use an nuclear bomb where a pea shooter will do?

OpenStudy (anonymous):

its ok you are learning. these mistakes must be made so they are not repeated in the future.

OpenStudy (anonymous):

Because explosions are cool, but I know what you mean

OpenStudy (tkhunny):

Fair enough. I'll give you that.

OpenStudy (anonymous):

Ok so I simplify as far as humanly possible first before I try to evaluate the limit?

OpenStudy (anonymous):

yes thats the idea

OpenStudy (anonymous):

if you were to plug in points you would see that it is approaching some value.

OpenStudy (tkhunny):

If you didn't spend enough time with Rational Functions before you got to the Calculus, you will have to go back and get a better background. Rational Functions are important. Never "plug in". Just resist.

OpenStudy (anonymous):

No, I understand. That was just me being a moron. Thanks guys!!!